OpenStudy (anonymous):

At their going-out-of-business sale, Bill's Bikes sold 42 bikes and had 58 left. Which equation could be used to find x, the number of bikes the store had originally? @kevindragonlord

OpenStudy (rhystic):

x - 42 = 58 would be your initial equation which means x = 58+42 -> x = 100.